A vending machine that takes only dimes and quarters contain 30 coins, with a total value of $4.20. How many of each coin are there?

Found 2 solutions by Labelx88, Alan3354:
Answer by Labelx88(1)

you have 8 quarters and 22 dimes. i didnt use an equation i just guessed 8 quartes then did the math for $2.20 worth of dime which is 22.
8 quarters=$2.00
22 dimes=+$2.20
if you need an equation as well let me know ill try and put one together.

Answer by Alan3354(69443)

A vending machine that takes only dimes and quarters contain 30 coins, with a total value of $4.20. How many of each coin are there?
---------------------
d + q = 30 (# of coins)
10d + 25q = 420 (total cents)
Can you do the rest?
====================
Another method:
30 dimes = $3.00
For each dime replaced by a quarter, 15 cents is added
420 - 300 = 120
120/15 = 8 quarters
22 dimes